OpenStudy (babyslapmafro):

Please help me solve the given inverse Laplace transform...(click to see)

OpenStudy (babyslapmafro):

\[L ^{-1} \frac{ s }{ (s ^{2}+4)(s+2) }\]

zepdrix (zepdrix):

Remember how to set up the Partial Practions? \[\Large \frac{s}{(s^2+4)(s+2)}\quad=\quad \frac{As+B}{s^2+4}+\frac{C}{s+2}\]

OpenStudy (babyslapmafro):

Ok I think I remember doing that

OpenStudy (babyslapmafro):

I know decomposition but I forgot this particular method

OpenStudy (babyslapmafro):

I'll try that